OpenStudy (anonymous):

Solve the equation: 2/5 - 3/x = 1/(5x)

OpenStudy (anonymous):

\[\frac{ 2 }{ 5 } - \frac{ 3 }{ x } = \frac{ 1 }{ 5x }\] what's the LCD?

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 } +\frac{ 7 }{ x } + \frac{ 32 }{ 6x }\] the LCD would be 6x for all the denominators can multiply into 6x 3 * 2x | x * 6 | 6x * 1 |

OpenStudy (anonymous):

what ever you multiply the denominator by you HAVE to multiply the same number to the numerator.

OpenStudy (anonymous):

x=8 ?

OpenStudy (anonymous):

let's go back to the original equation and look at that in the denominator what is the highest number they can all equal? 5, x, and 5x

OpenStudy (anonymous):

lowest number*

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

touche -.-" lol i meant without dividing 5, x and 5x can ALL equal 5x

OpenStudy (anonymous):

I have no idea ;'(

jimthompson5910 (jim_thompson5910):

2/5 is equivalent to (2x)/(5x) if you multiply top and bottom by x -3/x is equivalent to 15/(5x) if you multiply top and bottom by 5 Notice how they now have the same denominator of 5x This means 2/5 - 3/x = 1/(5x) is the same as (2x)/(5x) - 15/(5x) = 1/(5x) which allows you to combine the two fractions on the left side to get (2x - 15)/(5x) = 1/(5x) What's next?

jimthompson5910 (jim_thompson5910):

any ideas?

OpenStudy (anonymous):

aaahhhhhh

jimthompson5910 (jim_thompson5910):

yes? no?

jimthompson5910 (jim_thompson5910):

If two fractions are equal, and their denominators are equal, then the numerators must be equal

jimthompson5910 (jim_thompson5910):

so this means that if (2x - 15)/(5x) = 1/(5x) then 2x-15 = 1

jimthompson5910 (jim_thompson5910):

if 2x - 15 = 1, then x = ???

OpenStudy (anonymous):

8?

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

Thank youu

jimthompson5910 (jim_thompson5910):

yw